HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 108     Entry time: Thu Apr 3 03:48:40 2025
Author: JEL, ZN 
Category: Runs 
Subject: Screenshots 
The peaks in the analyzer are slowly shifting back to the original position. 
Note that a quadrupole was the cause of the issue 45 mins ago and it was cycled.  Might be relevant...
Attachment 1: 2025-04-03_03-47-28-530.png  362 kB  Uploaded Thu Apr 3 04:50:11 2025  | Hide | Hide all
2025-04-03_03-47-28-530.png
Attachment 2: 2025-04-03_03-47-11-337.png  390 kB  Uploaded Thu Apr 3 04:50:16 2025  | Hide | Hide all
2025-04-03_03-47-11-337.png
Attachment 3: 2025-04-03_03-47-42-665.png  640 kB  Uploaded Thu Apr 3 04:50:20 2025  | Hide | Hide all
2025-04-03_03-47-42-665.png
ELOG V3.1.5-fc6679b